If the line joining  (1, 2) and (4, 3) is divided by the line a + y = 5 then the division ratio is:
A. -1:1
B. 1:1
C. 1:2
D. 2:1

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
hello pushpa i think there is a slight mistake in ques as it should prolly be x+y= 5
now, we first find eqn of line joining (1, 2) and (4, 3).
y – 2= [(3 – 2)/(4 – 1)]*(x – 1)
3y – 6= x – 1
or 3y= x+5
now, we find intersection point of this line with x+y= 5.
we get it by solving the 2 eqns simultaneously: (5/2, 5/2).
now, if (5/2, 5/2) divides (1, 2) and (4, 3) in the ratio k:1, then
(4k+1)/(k+1)= 5/2
or k=1
so ratio is 1:1.
B is correct.
kindly approve :)